Use the Principle of Mathematical Induction to prove that (2𝑛)! /(2n)𝑛! is odd for all positive integers.


1
Expert's answer
2022-01-03T16:30:35-0500

ForΒ n=1, we have

"\frac{(2 n) !}{2 ! n !}=\frac{2 !}{2 ! !}=1 \text { is odd }"

Hence the result is true for n=1 (base case)

Let it be true for n=K.

Now for n=k+1

"\begin{aligned}

\frac{(2 k+2) !}{2^{k+1} \cdot(k+1) !} &=\frac{(2 k+2)(2 k+1)(2 x) !}{2 \cdot 2^{k} \cdot(k+1) \cdot k !} \\

&=\frac{(2 k) !}{2^{k} \cdot k !} \cdot \frac{2(k+1) \cdot(2 k+1)}{2(k+1)} \\

&=\frac{(2 k) !}{2^{k} \cdot k !} \cdot(2 k+1) \\
&\text{where}\ \frac{(2 k) !}{2^{k} \cdot k !} ,(2 k+1)\ \text{are odd}
\end{aligned}"

Hence the result is true for n=k+1.

Thus the result is true for all +ve integers by mathematical induction.